Message boards : News


log in

Advanced search
Message boards : News

Previous · 1 . . . 6 · 7 · 8 · 9 · 10 · Next
Threads Posts Author Views Last post
lockedWe fixed the bogus URL bug and all servers are back up
1 SETI News 3325 3625 days ago
lockedA trio of splitters are now up and producing new work. We plan to release major version 4 of BOINC and SETI@Home tomorrow. A clarification about last week's fallback to an earlier DB state: all work that was uploaded
1 SETI News 3555 3626 days ago
lockedThe validator is up and running and credit is now being calculated for all those backlogged results.
1 SETI News 3265 3628 days ago
lockedWe are currently working on getting the alpha/beta projects working again, as well as getting new workunits generated so that when we restart the public SETI@home project there will be work to send out to the clients. Another n
1 SETI News 4292 3631 days ago
lockedThe database has been restored and is being checked now
1 SETI News 3624 3632 days ago
lockedWe are now recovering the database from a snapshot made a few days ago
1 SETI News 3312 3633 days ago
lockedThe project ended up running all night despite the dropouts, but this afternoon we encountered a sudden raid failure that caused some data corruption
1 SETI News 3317 3634 days ago
lockedToday we are continuing to work with Snap technical support to solve our mounting problems
1 SETI News 3611 3635 days ago
lockedCurrently the project is up, as we were unable to recreate the problems we were seeing on Friday
1 SETI News 3425 3635 days ago
lockedWe encourage everyone to run SETI@home Classic during our server outages. SETI@home Classic makes the same scientific contribution as SETI@home/BOINC.
1 SETI News 3372 3637 days ago
lockedAll the data has been copied to the Snap device
1 SETI News 4007 3638 days ago
lockedThe download directory is fully transferred at this point, but we are still waiting for the upload directory to finish
1 SETI News 3601 3639 days ago
lockedThe DB is now on the Snap Appliance and the web site is back up. The project itself is still down as the transfer of the upload and download directories is still under way.
1 SETI News 3219 3639 days ago
lockedThe project is now down for the move of the DB and the upload directory to the Snap Appliance
1 SETI News 3282 3640 days ago
lockedWe have started the copy of the download directory to the Snap Appliance 18000.
1 SETI News 3409 3641 days ago
lockedThe servers are back up as far as receiving results but not as far as distributing new workunits
1 SETI News 3706 3643 days ago
lockedThe scheduling server is down while we work on a download storage problem
1 SETI News 3673 3644 days ago
lockedWe developed a hierarchical directory structure to get around the big directory problem (see August 4, below)
1 SETI News 4191 3645 days ago
lockedA misconfigured scheduling server ran for about an hour this morning, misinforming users that they had to upgrade to version 4 of the BOINC software
1 SETI News 3467 3646 days ago
lockedWe discovered that the sheer number of files in the upload directory was killing NFS performance for the file_upload_handler
1 SETI News 3939 3647 days ago
lockedWhile the project is up and functioning normally, the database and web servers are overloaded
1 SETI News 3260 3648 days ago
lockedThe project remained down all weekend, but we are back up now. To add insult to injury, the entire lab lost network connectivity to the rest of the world for a few hours this morning. In light of the unexpected length of this las
1 SETI News 3375 3649 days ago
lockedThe project is *still* down
1 SETI News 3514 3652 days ago
lockedThe project is currently down
1 SETI News 3276 3654 days ago
lockedThe project is down while we deal with a backup problem
1 SETI News 3372 3657 days ago
lockedWe have shutdown the schedulers and file upload handlers so the transitioners can play catch up again
1 SETI News 3354 3657 days ago
lockedThe good news is that the scheduler is now running as a fast cgi which means it maintains a persistent connection to the DB
1 SETI News 3631 3659 days ago
lockedWe are now up and running again, during the night we generated 400k results to be sent out
1 SETI News 3296 3659 days ago
lockedWe found and fixed the no results bug
1 SETI News 3388 3660 days ago
lockedYesterday we introduced a bunch of performance enhancements to the system
1 SETI News 3415 3660 days ago
lockedThe log files are currently being compressed, after thats done the project should return to normal.
1 SETI News 3620 3661 days ago
lockedWe have brought up an additional transitioner to try and cut through the backlog we have accrued
1 SETI News 3531 3663 days ago
lockedSometime between 5pm and 12am PDT, the splitters seem to have stopped producing new workunits
1 SETI News 3442 3663 days ago
lockedA server bug was preventing new results from being created in response to failed results. If 2 out of 3 results for a WU succeeded and the 3rd failed, the 2 successful results would not be validated or credited. This is no
1 SETI News 3409 3665 days ago
lockedOur database server hung and had to be forcibly terminated. It took several hours to check all the tables and indices. Then we ran into problems with our Network Attached Storage device, which we rebooted for the first tim
1 SETI News 3736 3669 days ago
lockedIt looks as though the server components (splitter, transitioner, feeder, scheduler) are now keeping up with demand.
1 SETI News 3482 3673 days ago
lockedSince yesterday evening we have distributed over 600,000 results
1 SETI News 3763 3674 days ago
lockedWe have moved the scheduling server to another machine in order to free up CPU and memory for the DB server
1 SETI News 3252 3674 days ago
lockedWe're making some major changes to deal with server performance problems. The project will be shut down for several hours today.
1 SETI News 3593 3675 days ago
lockedWe made some progress and are turning the project on. You may still have problems getting work. If that is the case, you might consider switching back to SETI@home Classic for the next few days.
1 SETI News 3688 3680 days ago
lockedWe've been having problems with server performance. To address these problems we're making some changes to our database. The project will be shut down most of today while this is being done.
1 SETI News 3972 3680 days ago
lockedYesterday morning we shut down the front end of the system to let the transitioners catch up with several days worth of backlogged work, which is why some of you were noticing three results and no credit assignment
1 SETI News 3755 3682 days ago
lockedAn additional transitioner has been brought online. We are turning on the web site and will be monitoring the system. We may have to shut it down again if traffic becomes too bad. We are investigating further performance enhanceme
1 SETI News 3482 3685 days ago
lockedThe web site has been reenabled. We have several upgrades scheduled throughout the week which should help with the performance problems. If the project runs out of work again you are welcome to revert back to the original
1 SETI News 3197 3685 days ago
lockedWe passed the safe threshold again, so the web site has been shut down until Monday, after we have had a chance to relocate some more infrastructure services.
1 SETI News 3421 3685 days ago
lockedWe are going to leave the project down until Monday. We are going to bring back up the scheduler in an hour. The transitioner was moved to its own machine and is processing workunits at a much faster rate.
1 SETI News 3584 3686 days ago
lockedMemory for our web server was defective; we are waiting on replacements before bringing it online.
1 SETI News 3551 3687 days ago
lockedWe added a second splitter and then hit the next bottleneck which was the transitioner.
1 SETI News 3350 3687 days ago
lockedThe transitioner is not currently able to keep up with the splitters. We are going to top the project to allow the transitioner a chance to catch up with the rest of the system.
1 SETI News 3576 3687 days ago
lockedTo handle increased demand for work, we are bringing another splitter online.
1 SETI News 3390 3688 days ago

Previous · 1 . . . 6 · 7 · 8 · 9 · 10 · Next

This message board is available as an RSS feed

Copyright © 2014 University of California